OpenStudy (anonymous):

Complete the square: i) x^4 +2x^2 +y^4 - 2y^2 +3

OpenStudy (anonymous):

The answer is supposed to be 2

ganeshie8 (ganeshie8):

for simplicity, say x^2 = m, y^2 = n

OpenStudy (yttrium):

Do you have any idea on applications of completing the squares?

ganeshie8 (ganeshie8):

x^4 +2x^2 +y^4 - 2y^2 +3 m^2 + 2m + n^2 - 2n + 3

OpenStudy (anonymous):

Applications? Not really. I've only ever answered questions with max powers of 2. I think ganeshie8's solution looks like it might work

OpenStudy (yttrium):

Yes. That will really work! Haha. Do you know that next steps?

OpenStudy (anonymous):

I've got (m+1)^2-1 + (n-1)^2 -1 +3. So then (m+1)^2 + (n-1)^2 +1

ganeshie8 (ganeshie8):

good job !

ganeshie8 (ganeshie8):

The answer is supposed to be 2 ??? dint get this part

OpenStudy (anonymous):

I then have to find the smallest value apparently (for real x and y)

ganeshie8 (ganeshie8):

you wanto find smallest value of function ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

let me think lol it looks tough for me :|

OpenStudy (anonymous):

Ok, yeh it is tough

ganeshie8 (ganeshie8):

@experimentX @hartnn

ganeshie8 (ganeshie8):

question is : find minimum value of x^4 +2x^2 +y^4 - 2y^2 +3

OpenStudy (experimentx):

did you complete the square?

OpenStudy (anonymous):

Yes, we got (m+1)^2 + (n-1)^2 +1 where m=x^2 and n=y^2

OpenStudy (experimentx):

there are two parts of the problem. x^4 +2x^2 +y^4 - 2y^2 +3 --------- this part's is always positive. so minimum value is zero for x=0 the toehr part y^4 - 2y^2 +3 complete square (y^2 - 1)^2 + 2 the you get the minimum value by minimizing (y^2 - 1)^2 ... that is y=1

OpenStudy (experimentx):

so the minimum value you get is when x=0 and y=1, that is 2

OpenStudy (anonymous):

Ah, I get it! Thank you!

ganeshie8 (ganeshie8):

thanks @experimentX it was awesome !